OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (anonymous):

subsititute the value of y=x+7 into the first equation x+7= -2x +19

OpenStudy (anonymous):

A. x = 4, y = 11 B. x = -12, y = -5 C. x = -4, y = 3 D. x = 8, y = 3 What is the answer of this a, b, c or d

OpenStudy (anonymous):

you have to work it out yourself too..

OpenStudy (anonymous):

Yes but if i'm here because I DON'T Know the ASWER :)

hero (hero):

(a) Set y = y: y = y (b) Set the expressions for y equal to each other: x+ 7 = -2x + 19 (c) Add 2x to both sides: 2x + x + 7 = 2x - 2x + 19 (d) Simplify 3x + 7 = 19 (e) Subtract 7 from both sides: 3x + 7 - 7 = 19 - 7 (f) Simplify 3x = 12 (g) Divide both sides by 3: 3x/3 = 12/3 (i) Simplify x = 4

hero (hero):

(j) substitute the value of x back into one of the original equations to find y: y = x + 7, x = 4, therefore: y = 4 + 7 y = 11 Therefore: x = 4, y = 11